OpenStudy (anonymous):

One frozen yogurt store sells frozen yougurt for $3.00 per cup and $1.25 per topping. Write a linear equation to show the total cost of a cup of frozen yogurt. Then calculate the total price for one cup of frozen yogurt with four toppings.

OpenStudy (anonymous):

google it

OpenStudy (anonymous):

y=3+1.25(x) y=3+1.25(4) y=3+5 y=8 $8.00

OpenStudy (anonymous):

Okay. Would this equation work? y=5x+3; $8.00

OpenStudy (anonymous):

No, I'm sorry. It has to be 1.25. You are not paying 5.00 a topping

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

yup:)

OpenStudy (anonymous):

Estimate the answer to 44,596,251 – 28,257,934 by first rounding each number to the nearest million.